OpenStudy (anonymous):

Explain, in complete sentences, the process of using Cramer’s Rule to solve the following system of equations. Provide the solution to the system. 2x + 2y = 14 x – 3y = –17

OpenStudy (anonymous):

Well first you do a determinant of the original which would be \[\left| \frac{ 2 }{ 1 }\frac{ 2 }{3 } \right|\]=(-6)-2=-8 Then replace the answer column with the x column and redo the determinant \[\left| \frac{ 14 }{ -17 }\frac{ 2 }{ -3 } \right|\]=(-42)-(-19)=-23 Put this value of the original and you got your x \[x=\frac{ 23 }{ 8 }\] Do same for y and you get the y value